# Towards Infant Sleep-Optimized Driving: Synergizing Wearable and Vehicle Sensing in Intelligent Cruise Control

### v2

@@ Abstract
Automated driving (AD) has substantially improved vehicle safety and driving comfort, but their impact on passenger well-being, particularly infant sleep, is not sufficiently studied. Sudden acceleration, abrupt braking, and sharp maneuvers can disrupt infant sleep, compromising both passenger comfort and parental convenience. To solve this problem, this paper explores the integration of reinforcement learning (RL) within AD to personalize driving behavior and optimally balance occupant comfort and travel efficiency. In particular, we propose an intelligent cruise control framework that adapts to varying driving conditions to enhance infant sleep quality by effectively synergizing wearable sensing and vehicle data. Long short-term memory (LSTM) and transformer-based neural networks are integrated with RL to model the relationship between driving behavior and infant sleep quality under diverse traffic and road conditions. Based on the sleep quality indicators from the wearable sensors, driving action data from vehicle controllers, and map data from map applications, the model dynamically computes the optimal driving aggressiveness level, which is subsequently translated into specific AD control strategies, e.g., the magnitude and frequency of acceleration, lane change, and overtaking. Simulation results demonstrate that the proposed solution significantly improves infant sleep quality compared to baseline methods, while preserving desirable travel efficiency.
@@ III-D Reward Function
In our work, reward is designed to capture user’s requirement on travel time. The reward for taking an action determined at the $i^{th}$ checkpoint is calculated at the next checkpoint, i.e., the $(i+1)^{th}$ checkpoint. In particular, the estimated time of arrival (ETA) is denoted as the $t^{E}_{i}$, which indicates the expected time needed to arrival at $i^{th}$ checkpoint based on the route planning result. Similarly, the actual time of arrival (ATA) is denoted as $t^{A}_{i}$, which is the actual time elapsed upon arriving at the $i^{th}$ checkpoint since the beginning of the trip. Then the reward for taking the action can be calculated by a reward function, written as
@@ III-F Termination Condition of Each Episode
Since the objective of the proposed approach is to improve infant sleep quality, each training episode should terminate when the baby wakes up. Multiple algorithms[17] have been developed to precisely estimate sleep state based on various sensors input. In this paper, we use estimate sleep state from motion sensors, e.g., IMU, since the sensor is broadly available in wearable devices. But more sophisticated algorithms using various sensor inputs, e.g., heart rate and respiratory rate, can be considered similarly. In particular, a large motion value detected by the motion sensor can indicate that the user has awakened, since the motion value is usually at a minimum level when the user is asleep. Therefore, the episode ends when the maximum motion $M_{i}^{\max}$ exceeds a threshold $M^{\text{thres}}$, i.e., when $M_{i}^{\max}\geq M^{\text{thres}}$.
@@ IV Network modules and algorithms
The Proximal Policy Optimization (PPO) algorithm [18] is adopted to train the proposed models. Given that the state vector encapsulates multidimensional temporal aspects of the driving environment, we propose a hybrid network architecture for effective processing. as shown in Fig. 3.
To achieve multi-dimensional information fusion and feature extraction, we employ a Multilayer Perceptron (MLP) network. For temporal feature learning, we investigate two deep neural network architectures: Long Short-Term Memory (LSTM) and Transformer models. These architectures are designed to effectively capture sequential dependencies and long-range temporal correlations inherent in diverse driving scenarios.
@@
where ${V}_{i}^{l}$ is output of the $l^{\text{th}}$ layer of MLP and $\bm{W}_{p}$, $\bm{b}_{p}$ are the corresponding weight and bias parameters. Then, the outputs $\bm{V}^{L}_{i-K+t}$, for $t=1,2,...,K$, from the last MLP layer is fed into the LSTM module, given by
$c_{t},h_{t}=\text{LSTM}(c_{t-1},h_{t-1}).$ (6)
The final hidden state $c_{t}$ derived from the last LSTM output is used to calculate final result of the network:
@@
where $\bm{Q}_{C}$, $\bm{K}_{C}$, and $\bm{V}_{C}$ are given by:
$\bm{Q}_{C}=\bm{C}_{t}\bm{W}_{Q},\quad\bm{K}_{C}=\bm{C}_{t}\bm{W}_{K},\quad\bm{V}_{C}=\bm{C}_{t}\bm{W}_{V}.$ (10)
Here, $\bm{Q}_{C}$, $\bm{K}_{C}$, and $\bm{V}_{C}$ represent the query, key, and value matrices derived from the input $\bm{C}_{t}$ using the learnable projection weights $\bm{W}_{Q}$, $\bm{W}_{K}$, and $\bm{W}_{V}$, respectively. The resulting attention output $\bm{Z}_{t}$ encodes temporal interactions within the sequence. The final output representation $\bm{O}_{i}$ of the environment state is the first token in $\bm{Z}_{t}$.
@@ V-A Simulation Setup
$a_{\text{wrist}}(t)=a_{\text{car}}(t)+\frac{1}{m}\left(-kx(t)-cv(t)\right),$ (13)
[…]
The PPO model outputs a continuous-valued driving aggressiveness level and is trained over 2000 iterations with 20 steps per update, using 100 different traffic condition settings and running routes in Town07. The Adam optimizer is used with a learning rate of $1\times 10^{-4}$, $\gamma=0.99$. Two policy architectures are evaluated: (1) an LSTM-based policy with sequence length 5, hidden size 128, 2 layers, and 0.1 dropout; and (2) a Transformer-based policy with 2 encoder layers, 4 attention heads, 128-dimensional embeddings, and learned positional encoding.
@@ V-B Result Discussion
Figure 4: Convergence process of the proposed algorithm for LSTM and Transformer based network structures
@@
Figure 5: Relative maximum movement of occupant vs late rate of trip with different algorithms.
Next we investigate the proposed approach in terms of sleep quality improvement and trip delay. Sleep quality is quantified by the relative maximum movement of the occupant, defined as the normalized peak value of the motion sensor (i.e., IMU) output during a trip. Trip delay is measured by the trip late rate, calculated as $\max\left({\left(t_{I}^{A}-t_{I}^{E}\right)}/{t_{I}^{E},0}\;\right)$, where $I$ denotes the final checkpoint of the trip, $t_{I}^{E}$ the ETA, and $t_{I}^{A}$ the ATA. As shown in Fig. 5, we calculate these metrics across 10 randomly sampled trip sequences for the proposed LSTM and Transformer (TF) algorithms, as well as three baselines: Actn8, Actn2, and Rand. In Actn8 and Actn2, a fixed aggressiveness level of 8 and 2, respectively, is used at each step. In Rand, the aggressiveness level is randomly sampled from $\mathcal{A}$. To visualize performance differences, samples are categorized based on metric thresholds. A sample is labeled good (large bubble) if both metrics are low, i.e., relative maximum movement $<0.4$ and trip late rate $<1$. It is labeled poor (small bubble) if either metric is high, i.e., relative maximum movement $>0.8$ or trip late rate $>2$. All other samples are classified as acceptable (medium bubble).
It can be seen that LSTM achieves the optimal trade-off between sleep quality and trip latency across most trip samples. While TF performs similarly in latency, it shows higher maximum movement in several cases (e.g., more samples with maximum movement $>0.6$), indicating lower sleep quality. Both proposed models outperform benchmarks: Actn8 shows more samples with maximum movement $>0.9$, and Actn2 exhibits more samples with late rate $>1.5$. These results confirm the design assumption that increased aggressiveness reduces latency but degrades sleep quality. The Rand, with randomly assigned aggressiveness, shows no consistent performance pattern.
@@ VI Conclusions
This work proposes an intelligent cruise control framework that integrates wearable sensing and vehicle data to enhance occupant comfort, with a focus on infant sleep, under varying driving conditions. Novel network modules are designed to model the impact of driving behavior on sleep quality and travel efficiency across diverse road and traffic scenarios. The network is trained using PPO-based reinforcement learning and outputs driving–aggressiveness control decisions by leveraging multimodal inputs, including sleep‐quality data from wearable sensors, action data from vehicle controllers, and route information from map applications. Simulation results demonstrate that the proposed approach effectively adapts to dynamic driving conditions and outperforms benchmark methods in optimizing sleep quality while maintaining desirable trip efficiency.

### v3

@@ Abstract
Automated driving (AD) has substantially improved vehicle safety and driving comfort, but their impact on passenger well-being, particularly infant sleep, is not sufficiently studied. Sudden acceleration, abrupt braking, and sharp maneuvers can disrupt infant sleep, compromising both passenger comfort and parental convenience. To solve this problem, this paper explores the integration of reinforcement learning (RL) within AD to personalize driving behavior and optimally balance occupant comfort and travel efficiency. In particular, we propose an intelligent cruise control framework that adapts to varying driving conditions to enhance infant sleep quality by effectively synergizing wearable sensing and vehicle data. Long short-term memory (LSTM) and transformer-based neural networks are integrated with RL to model the relationship between driving behavior and infant sleep quality under diverse traffic and road conditions. Based on the sleep quality indicators from the wearable sensors, driving action data from vehicle controllers, and map data from map applications, the model dynamically computes the optimal driving aggressiveness level, which is subsequently translated into specific AD control strategies, e.g., the magnitude and frequency of acceleration, lane change, and overtaking. Simulation experiments conducted in the CARLA environment indicate that the proposed solution significantly improves infant sleep quality compared to baseline methods, while preserving desirable travel efficiency.
@@ III-D Reward Function
In our work, reward is designed to capture user’s requirement on travel time. Specifically, we compare the actual time of arrival (ATA) and the estimated time of arrival (ETA) at each checkpoint to assess how well the system adheres to the expected schedule. Let $t^{E}_{i}$ denote the ETA at the $i^{\text{th}}$ checkpoint, based on the route planning result, and $t^{A}_{i}$ denote the ATA, i.e., the actual elapsed time from the beginning of the trip to the $i^{\text{th}}$ checkpoint. The reward for taking an action at the $i^{\text{th}}$ checkpoint is determined at the $(i+1)^{\text{th}}$ checkpoint, and is defined by
@@ III-F Termination Condition of Each Episode
Since the objective of the proposed approach is to improve infant sleep quality, each training episode should terminate when the baby wakes up. Multiple algorithms[17] have been developed to precisely estimate sleep state based on various sensors input. In this paper, we estimate sleep state using motion sensors, e.g., IMU, as they are widely available in wearable devices. But more sophisticated algorithms using additional inputs, e.g., heart rate and respiratory rate, can be considered similarly. In particular, a large motion value detected by the motion sensor can indicate that the user has awakened, since the motion value is usually at a minimum level when the user is asleep. Therefore, the episode ends when the maximum motion $M_{i}^{\max}$ exceeds a threshold $M^{\text{thres}}$, i.e., when $M_{i}^{\max}\geq M^{\text{thres}}$.
@@ IV Network modules and algorithms
The Proximal Policy Optimization (PPO) algorithm [18] is adopted to train the proposed models. Given that the state vector encapsulates multi-dimensional temporal aspects of the driving environment, we propose a hybrid network architecture for effective processing. as shown in Fig. 3.
To achieve multi-dimensional information fusion and feature extraction, we employ a Multilayer Perceptron (MLP) network. For temporal feature learning, we investigate two deep neural network architectures: Long Short-Term Memory (LSTM) and Transformer models, as illustrated in Fig. 3. These architectures are designed to effectively capture sequential dependencies and long-range temporal correlations inherent in diverse driving scenarios.
@@
where $\bm{V}_{i}^{l}$ is output of the $l^{\text{th}}$ layer of MLP and $\bm{W}_{p}$, $\bm{b}_{p}$ are the corresponding weight and bias parameters. Then, the outputs $\bm{V}^{L}_{i-K+t}$, for t=1,2,,Kt=1,2,...,K, from the last MLP layer is fed into the LSTM module, given by
$\bm{c}_{t},\bm{h}_{t}=\text{LSTM}(\bm{c}_{t-1},\bm{h}_{t-1}).$ (6)
The final hidden state $\bm{c}_{t}$ derived from the last LSTM output is used to calculate final result of the network:
@@
where $\bm{Q}_{C}=\bm{V}_{t}\bm{W}_{Q}$, $\bm{K}_{C}=\bm{V}_{t}\bm{W}_{K}$, and $\bm{V}_{C}=\bm{V}_{t}\bm{W}_{V}$, and they represent the query, key, and value matrices derived from the input $\bm{V}_{t}$ using the learnable weights $\bm{W}_{Q}$, $\bm{W}_{K}$, and $\bm{W}_{V}$, respectively. The attention output $\bm{Z}_{t}$ encodes temporal interactions within the sequence. The final output representation $\bm{O}_{i}^{final}$ of the environment state is the first token in $\bm{Z}_{t}$.
@@ V-A Simulation Setup
$\alpha_{\text{wrist}}(t)=\alpha_{\text{car}}(t)+\frac{1}{m}\left(-kx(t)-cv(t)\right),$ (12)
where $\alpha_{\text{car}}$ is the vehicle acceleration retrived from onboard IMU and $v(t)$ is the vehicle speed. The configuration parameters are set as $m=0.4\text{kg}$, $k=50\text{N/m}$, and $c=1.0\text{Ns/m}$. The IMU magnitude is monitored over a 3-second (300-sample) window, and the infant is labeled awake if $\max\lVert\alpha_{\text{wrist}}\rVert>2\,\text{m/s}^{2}$. Aggressiveness level $a_{i}$ is selected from the space $\mathcal{A}=\{0,1,...,10\}$ and is implemented in Autopilot through configuration parameters, including Speed difference percentage $P_{s}$, Following distance $P_{\text{d}}$ (m), Path update interval $P_{\text{u}}$ (ms), Auto Lane change enable $P_{\text{a}}$, and Keep right percentage $P_{\text{r}}$. These parameters are computed as follows: $P_{\text{s}}=1.5\times a_{i}$, $P_{\text{d}}=6.0-0.4\times a_{i}$, $P_{u}=1000-70\times a_{i}$, $P_{\text{a}}=\mathbb{I}\{a_{i}\geq 2\}$, and $P_{r}=100-10\times a_{i}$.
The PPO model outputs a continuous-valued driving aggressiveness level and is trained over 2000 iterations with 20 steps per update, using 100 different traffic condition settings and running routes in Town07. The Adam optimizer is used with a learning rate of $1\times 10^{-4}$, $\gamma=0.99$. Two policy architectures are evaluated: (1) an LSTM-based policy with sequence length 5, hidden size 128, 2 layers, and 0.1 dropout; and (2) a Transformer-based policy with 2 encoder layers, 4 attention heads, 128-dimensional embeddings, and learned positional encoding. Training the RL agent takes approximately 30 minutes on a single NVIDIA RTX 3060 GPU.
@@ V-B Result Discussion
Figure 4: Convergence process of the proposed algorithm for LSTM and Transformer based network structures.
@@
Figure 5: Relative maximum movement of occupant vs late rate of trip with different algorithms (both axes shown as ratios).
Next we investigate the proposed approach in terms of sleep quality improvement and trip delay. Sleep quality is quantified by the relative maximum movement of the occupant, defined as the normalized peak value of the motion sensor (i.e., IMU) output during a trip. Trip delay is measured by the trip late rate, calculated as $\max\left({\left(t_{I}^{A}-t_{I}^{E}\right)}/{t_{I}^{E},0}\;\right)$, where $I$ denotes the final checkpoint of the trip, $t_{I}^{E}$ the ETA, and $t_{I}^{A}$ the ATA. As shown in Fig. 5, we calculate these metrics across 10 randomly sampled trip sequences for the proposed LSTM and Transformer (TF) algorithms, as well as three baselines: Actn8, Actn2, and Rand. In Actn8 and Actn2, a fixed aggressiveness level of 8 and 2, respectively, is used at each step. In Rand, the aggressiveness level is randomly sampled from $\mathcal{A}$. To visualize performance differences, samples are categorized based on metric thresholds. A sample is labeled good (large bubble) if both metrics are low, i.e., relative maximum movement <0.4<0.4 and trip late rate <1<1. It is labeled poor (small bubble) if either metric is high, i.e., relative maximum movement >0.8>0.8 or trip late rate >2>2. All other samples are classified as acceptable (medium bubble).
It can be seen that LSTM achieves the optimal trade-off between sleep quality and trip latency across most trip samples. While TF performs similarly in latency, it shows higher maximum movement in several cases (e.g., more samples with maximum movement >0.6>0.6), indicating lower sleep quality. Both proposed models outperform benchmarks: Actn8 shows more samples with maximum movement >0.9>0.9, and Actn2 exhibits more samples with late rate >1.5>1.5. These results confirm the design assumption that increased aggressiveness reduces latency but degrades sleep quality. The Rand, with randomly assigned aggressiveness, shows no consistent performance pattern.
@@ VI Conclusions
This work presents an intelligent cruise control framework that combines wearable sensing and vehicle data to enhance occupant comfort, with a focus on infant sleep under varying driving conditions. The system leverages reinforcement learning with novel network modules to model the impact of driving behavior on sleep quality and travel efficiency. It generates driving aggressiveness decisions based on multimodal inputs, including sleep-related signals from wearables, vehicle control data, and route information. Simulations show that the approach adapts effectively to dynamic driving conditions and outperforms baselines in improving sleep quality without sacrificing trip efficiency. Future work will address practical deployment challenges such as sensor latency and system integration, and explore additional physiological signals for more accurate sleep state estimation.